OpenStudy (anonymous):

0.002 = 0.003 - 0.002 / 1 - F^2 How do i rearrange this for F?

OpenStudy (akashdeepdeb):

Is this the question: \[0.002 = \frac{0.003-0.002}{1-F^2}\] Or \[0.002 = 0.003 - \frac{0.002}{1-F^2}\] ?

OpenStudy (anonymous):

the answer is 0.71, i get 1.41??

OpenStudy (anonymous):

@AkashdeepDeb it is the top one

OpenStudy (akashdeepdeb):

Okay, so your question is: \[0.002 = \frac{0.003-0.002}{1-F^2}\] Now you can simplify the numerator on the RHS [Right Hand Side] 0.003 - 0.002 = 0.001 You'd get: \[0.002 = \frac{0.001}{1-F^2}\] Now multiply both sides by 1000 [The decimals places look confusing :P ] You'd get: \[2 = \frac{1}{1-F^2}\] Now re-arrange the equation, by taking \(1-F^2\) to the LHS. And bringing 2 down. Then you'd get: \[1-F^2 = \frac{1}{2} = 0.5\] Subtract 1 from both sides then. And then multiply both sides by -1. You'd get the value of F after doing that. :D Did you get this? :)

OpenStudy (anonymous):

do u get 0.71?

OpenStudy (akashdeepdeb):

I think you should try that out. But you'd require a calculator. If you understood, all what I did there. You should, most probably, get the answer. Shall I explain again?

OpenStudy (anonymous):

thanks bra

OpenStudy (anonymous):

how is the ipl going?

OpenStudy (anonymous):

in your last step why does the 2 come down? when u bring it accross

OpenStudy (akashdeepdeb):

You divide both sides by 2. And multiply both sides by 1-F^2

OpenStudy (anonymous):

isnt the 2 just a positive number? u bring it accross as a minus

OpenStudy (akashdeepdeb):

We found that: \[2 = \frac{1}{1-F^2}\] Now what we do is, we divide both sides by 2. [As it is an EQUALITY we have to do the same things on both the sides, so we divide BOTH the sides by 2] The main aim, is to get the value of 'F'. So we need to get 'F' to one of the sides of the equality to calculate its value. We get: \[\frac{2}{2} = 1 = \frac{1}{2(1-F^2)}\] Then we multiply both sides by \(1-F^2\) To get the 'F' on one of those sides. Then you'd get: \[1-F^2 = \frac{1}{2}\] And then you just rearrange the equation to get the value of 'F'. Did you get this? :)

OpenStudy (anonymous):

thankyou i truly appreciate your great explination...

OpenStudy (akashdeepdeb):

You're welcome. :)